OpenStudy (karatechopper):

difference between (-5)^2 and -5^2?

OpenStudy (compassionate):

There isnone.

OpenStudy (anonymous):

Are you sure there is a diff. Cause if there is then it has to be the fact that (-5)^2 has the negative in it too so the negative sign gets squared too

OpenStudy (anonymous):

Making it positive

OpenStudy (karatechopper):

well on my worksheet, there is a problem for those two. And i thought that (-5)^2 would be -5 x -5=25 and that -5^2= -1(5^2)=-1(25)=-25

OpenStudy (anonymous):